eBook Bank A pays 7% interest compounded annually on deposits, while Bank B pays 6.5% compounded...

eBook

Bank A pays 7% interest compounded annually on deposits, while Bank B pays 6.5% compounded daily.

a. Based on the EAR (or EFF%), which bank should you use?

b. Could your choice of banks be influenced by the fact that you might want to withdraw your funds during the year as opposed to at the end of the year? Assume that your funds must be left on deposit during an entire compounding period in order to receive any interest
Answer #1

a. EAR for Bank A is 7% because compounding is annual and 7% interest rate is also annual.

EAR for Bank B = (1+interest rate/no. of compounding)no. of compounding - 1

in daily compounding, no. of compounding will be 365 as there are 365 days in a year.

EAR for Bank B = (1+0.065/365)365 - 1 = 1.000178082191781365 - 1 = 1.0672 - 1 = 0.0672 or 6.72%

You should use Bank A because based on EAR bank A pays higher interest at 7% rate.

b. your choice of banks could be influenced by the fact that you might want to withdraw your funds during the year as opposed to at the end of the year if your funds must be left on deposit during an entire compounding period in order to receive any interest. in this case your choice should be Bank B because it has daily compounding. Bank A will pay interest only once in a year and that also at the end of year. so if you withdraw your funds during the year then Bank A will not pay any interest to you.